HUBBLE SPACE TELESCOPE DAILY REPORT #2271 PERIOD COVERED: 0000Z (UTC) 12/15/98 - 0000Z (UTC) 12/16/98 Daily Status Report as of 350/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell Nova Sgr-1 in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.3 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was to be used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. Due to the NICMOS suspend event that occurred on day 348, this proposal did not execute. 1.4 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Three Sets of WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxies FCC212, VCC646 and DGC78 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Eleven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. Due to the NICMOS suspend event that occurred on day 348, all but the final iteration of this proposal, which completed with no reported errors, did not execute. 1.8 Completed Three Sets of WF/PC-2 6751 (Is the Central Star of K 648 in M15 a Close Binary?) The WF/PC-2 was used to make observations of the central star of the planetary nebula K648 which is located in globular cluster M-15 to determine if this star is a binary star. Pre-repair HST images reveal that the nebula is strongly elliptical in shape, and shows blobs on either side of the central star. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Seven Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 7358 (The Unusual Absorption Line System of PG2302+029 -- Ejected or Intervening?) The WF/PC-2 was used to make observations of the unusual absorption line systems PG2302+029 to help to identify the cause of this absorption system and allow us to determine whether the system is an extreme example of previously known classes of quasar absorption lines or represents an entirely new phenomenon. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxy NGC-5077 (taken form a list of 150 galaxies) in order to study its nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed FGS 7678 (Monitor the Long-Term Stability of FGS-1R) The goal of these measurements is to monitor the internal positional stability of the FGSs by measuring many stars in the M35 check field repeatedly over a one year period. The results of these measurements are the stability characteristics of the FGS to allow an estimate of differential corrections to OFAD and Plate Scale. The observations were completed as planned, and no problems were reported. 1.13 Completed STIS/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (MA2) was used to make observations of the star BD+75D325 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) PK242-11D1 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena, yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.15 Completed Five Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel Low galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 15 Successful : 15 Scheduled Reacquisitions : 2 Successful : 2 2.2 FHST Updates: Scheduled : 34 Successful : 31 As documented in HSTAR 6817, the full maneuver updates at 349/195249Z and 349/195534Z failed. The following acquisition was successful in spite of a tracker update failure immediately prior. A map taken after the acquisition showed small errors. 2.3 Operations Notes: The NICMOS instrument began the day in a suspend state. A meeting held early Tuesday morning found no indications of a hardware failure. Memory dump analysis showed no apparent problems. Consensus was to recover the instrument as soon as practical. Accordingly, RTCS 25 was managed at the 349/1929Z, and the NICMOS was recovered from suspend via direction from an operations request. The instrument is now operating nominally. The proposals detailed in 1.3 and 1.7 were affected by the suspend. Using ROP SR-1, the SSR EDAC error counter was cleared six times. The STIS MAMA-1 MCE reset at 349/184854Z during an SAA event and while the low voltage was on. Per ROP NS-12, the STIS error counter was reset at 349/1900Z. The MCE was recovered at 350/080648Z, the time of a MAMA turn-on, via SMS commanding. HSTAR 6816 was written. There was a NICMOS EDAC single bit error at 350/083655Z. HSTAR 6818 was written, and the ground system limit was adjusted. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH